OpenStudy (anonymous):

Check if I'm correct? (-8,5) and (-8,8) the slope is 3 or 0?

OpenStudy (skullpatrol):

Neither.

OpenStudy (skullpatrol):

Draw a sketch please.

OpenStudy (anonymous):

0

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):

What is the slope of the line passing through those two points?

OpenStudy (dls):

8

OpenStudy (dls):

tilt it

OpenStudy (skullpatrol):

@MeganEvans Are you still there?

OpenStudy (skullpatrol):

@MeganEvans Do you recall the definition of slope?

OpenStudy (dls):

wth?

OpenStudy (dls):

3/0= 3? good good @uri first learn some basic math urself

OpenStudy (dls):

what exactly is the "Abusive" language?

OpenStudy (psymon):

Rise over run \[\frac{ 3 }{ 0 }\] We cannot have a 0 in the denominator of a fraction, therefore undefined slope.

OpenStudy (skullpatrol):

$$\huge \frac{3}{0} \neq 3$$

OpenStudy (dls):

 \[\Huge m=slope=\tan90^0\]